The front 1.25 m of a 1385-kg car is designed as a ‘crumple zone’ that collapses to absorb the shock of a collision. If a car travelling 27.2 m/s stops uniformly in 1.28 m, a) how long does the collision last, b) what is the magnitude of the average force on the car, and c) what is the acceleration of the car? Express the acceleration as a multiple of the acceleration due to gravity.
Answer #1

(a) v²= u²+2 as T 0 = (27.2)² + 2x ax 1.28 - 289 => a = m/s² 27.2 - 289 € 0= & v= utat » t = 0.094 s Anower T (6) 1385 x 289